Federico and Giulia are two students who majored in foreign languages and who have recently

been hired by a company that translates documents. Each of them can spend their days

translating documents from English to Italian or Spanish. If Federico spends all his day

translating documents to Italian with no Spanish, he can translate 10 documents, and if

he spends all his day translating documents to Spanish with no Italian, he can translate 8

documents. If Giulia spends all her day translating documents to Italian with no Spanish, she

can translate 8 documents, and if she spends all her day translating documents to Spanish

with no Italian, she can translate 20 documents.

(a) The company requires its workers to translate 4 documents to Italian and 6 documents

to Spanish in a day. Is this feasible for Federico? How about Giulia?

(b) Federico and Julia agree to trade documents. Who should specialize in translating

documents to Italian if they trade? And who should specialize in translating documents

to Spanish if they trade?

(c) What is the lowest price of translating a document to Italian (in terms of documents

translated to Spanish) that will allow trade to occur? What is the highest price? (Hint:

the seller of Italian translations wants high prices and the buyer wants low prices, and

both must agree in order for trade to occur.)
Answer #1

A)if translate 4 Italian ,then 4/10 of the day will spend on the Italian and remaining 6/10=3/5=0.6

To translate 6 Spanish ,he requires 6/8 =3/4=0.75 but he has left only 0.6 of the day ,so it is not feasible for Federico

If giliua translate 4 Italian then, 4/8=1/2=0.5 of her day spend on Italian.she left with 0.5 of day

To translate 6 Spanish ,she needs 6/20 =3/10=0.3 but she has 0.5 of day left .

So it is feasible for guliliu .

B) Federico opportunity cost of translate Italian=8/10=0.8

Giliua opportunity cost of translate Italian=20/8=2.5

Federico has lower opportunity cost of translate Italian ,so he would specialize in translate Italian.

Federico opportunity cost of translate Spanish=10/8=1.25

Giliua opportunity cost of translate Spanish=8/20=0.4

Guliliu has lower opportunity cost of translate Spanish ,so she has comparitive advantage in translate Spanish,so she would specialize in translate Spanish.

C)the lowest price is opportunity cost of Federico as he has the lower opportunity cost=0.8

The highest price is opportunity cost of guliliu as she has higher opportunity cost of translate Italian=2.5

The seller of Italian { Federico} willing to sell Italian translate at price Equal to his opportunity cost=0.8

The buyer of Italian { giliua} willing to pay for Italian translate price equal to opportunity cost at which he can produce Italian=2.5